OpenStudy (anonymous):

The elimination process 2a+3b=4 -2a+b=12

OpenStudy (anonymous):

plus so 4b=16....then get b and a

OpenStudy (anonymous):

so how do you get the b and a

OpenStudy (anonymous):

2a + 3b = 4 -2a + b = 12 Just add linearly. 2a + 3b = 4 -2a + b = 12 ------------ 0 + 4b = 16 4b = 16 b = 4 Then plug 4 in for a. 2a + 3b = 4 2a + 3(4) = 4 Just solve for a now :)

OpenStudy (anonymous):

Thank you so very much

OpenStudy (anonymous):

Anytime :)